Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $120 per unit. Variable expenses are $60.006. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $1.80 per u6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $1.80 per u6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $1.80 per u6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $1.80 per u6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $1.80 per u6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $1.80 per u6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $1.80 per u6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $1.80 per u6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $1.80 per u
Answer #1

Solution 1:

CM ratio = Contribution margin / Sales = $1,620,000 / $3,240,000 = 50%

Solution 2:

Break even Point in dollar sales = Fixed costs / contribution margin ratio = $160,000 / 50% = $320,000

Solution 3:

If sales increases by $60,000, then increase in net operating income = $60,000 * 50% = $30,000

Solution 4a:

Degree of operating leverage based on last year sales = contribution margin / net operating income

= $1,620,000 / $1,460,000 = 1.11

Solution 4b:

% increase in net operating income = % increase in sales * Degree of operating leverage

= 14% * 1.11 = 15.54%

Solution 5a:

New selling price per unit = $120 * 86% = $103.20

New Fixed costs = $160,000 + $69,000 = $229,000

New sales volume = ($3,240,000/$120)*125% = 33750 units

Computation of net operating income - Feather friend Inc.

Particulars

Amount

Sales (33750 * $103.20)

$3,483,000.00

Variable expenses (33750*$60)

$2,025,000.00

Contribution margin

$1,458,000.00

Fixed expenses

$229,000.00

Net operating income

$1,229,000.00

Solution 5b:

If the sales manager's ideas are implemented, net increase (Decrease) in net operating income = $1,229,000 - $1,460,000 = ($231,000)

As net operating income is decreasing as compared with last year, therefore sales manager suggestions should not be implemented.

Solution 6:

If sales commission increased by $1.80 per unit then new contribution margin per unit = $120 - $60 - $1.80 = $58.20 per unit

New sales volume = 27000*125% = 33750 units

Target operating income = $1,460,000

Maximum fixed costs = Total contribution margin - target operating income

= (33750 * $58.20) - $1,460,000 = $504,250

Existing fixed costs = $160,000

The amount by which advertising can be increase is = $504,250 - $160,000 = $344,250